OpenStudy (anonymous):

A rectangular city pack has an area of 211,200 squARE FEET IF THE LENGTH OF THE PARK IS 660 FT. WHAT IS THE WIDTH OF THE PARK

OpenStudy (anonymous):

Imeant -27

OpenStudy (anonymous):

-27= 3/4 w Multiply by reciprocal (4/3) * -27 = (3/4)*(4/3)w this will cancel out the 3/4 and isolate the w now you have (4/3)*-27 = w then by reducing the 3 and -27 you get 4*(-9) = w therefore you get -36 = w Make sense?

OpenStudy (anonymous):

yep